# Large Language Model Enabled Multi-Task Physical Layer Network

### v1

@@ III-E Multi-task Loss Function and Training Schedule
During the training stage, we randomly choose a task and sample data from the corresponding dataset for each iteration. After determining the involved task and data, we select the proper prompt and activate the required modules of encoder and decoder. It is worth noting that the backbone of the pre-trained LLM is frozen, while the other parameters of the network, together with the LoRA adapters are trainable. Then the proposed network update the parameters of encoders, decoders and LoRA adapters using the corresponding loss. Next, we will illustrate the loss function of the tasks, respectively.
@@
For multi-user precoding, the training is divided into two stages, namely supervised learning and unsupervised learning, respectively. The WMMSE algorithm is one of the most popular iterative to find the locally optimal solutions. Therefore, in the supervised learning stage, we can obtain the power allocation vectors $\underline{\mathbf{p}}$ and $\underline{\mathbf{\lambda}}$ according to the WMMSE algorithm as the label. Then the supervised learning with the loss function based on the MSE metric will be first used to achieve as close to the results of the WMMSE algorithm as possible, i.e.,
@@
$\text{Loss}_{Pre}=\frac{1}{2K}\left(\|\underline{\mathbf{p}}-\hat{\mathbf{p}}^{*}\|_{2}^{2}+\|\underline{\bm{\lambda}}-\hat{\bm{\lambda}}^{*}\|_{2}^{2}\right),$ (26)
where $\underline{\mathbf{p}}$ and $\underline{\bm{\lambda}}$ are the power vectors obtained from the WMMSE algorithm. However, the WMMSE algorithm is locally optimal and thus (26) is not equivalent to the real objective of problem P3 which aims to maximize the weighted sum rate. To further improve the sum rate performance, we continue to train the network in an unsupervised learning way, whose loss function takes the objective function directly as a metric, i.e.,
@@
In the signal detection task, the original transmitted data is rearranged as $\mathbf{X}_{\rm Dec}\in\mathbb{R}^{K\times 2}$, which serves as ground truth of the network output. Then we choose MSE loss for the training phase,
$\text{Loss}_{Dec}=\frac{1}{2K}\|\mathbf{X}_{\rm Dec}-\mathbf{X}_{\rm Dec}^{O}\|_{F}^{2}.$ (28)
@@
IV Simulation Results
IV-A Simulation Setup
In the simulation parts, we adopt the widely used channel generator QuaDRiGa [25], where we consider the 3GPP Urban Macro (UMa) channel model [26] and non-line-of-sight (NLOS) scenarios. The number of clusters is 21 and the number of paths per cluster is 20. We consider a multi-user MISO-OFDM system, where a BS simultaneously serves $K=4$ moving users. BS is equipped with a UPA with $N_{h}=16$, $N_{v}=8$ while each user is equipped with a single omnidirectional antenna. The antenna spacing is half of the wavelength at the center frequency. The users are uniformly distributed within angle range $[-\pi/2,\pi/2]$, and distance range $[\rho_{\rm min},\rho_{\rm max}]=[20~{}{\rm m},100~{}{\rm m}]$. We suppose a time-division duplex (TDD) system, where the center frequency of channel is set as $2.4$ GHz. The bandwidth of the channel is 8.64 MHz, comprising of $M=48$ subcarriers, i.e., the frequency interval of subcarriers is $180$ kHz. The training dataset and test dataset respectively contain 90000 and 10000 samples for each task.
For channel prediction problem, We predict future CSI of $T_{2}=4$ time slots based on historical CSI of $T_{1}=16$ time slots and set the time interval of pilots as $0.5$ ms. The initial position of the user is randomized and the motion trajectory is set as linear type. The user velocities are uniformly distributed between $10$ and $100$ km/h. For uplink signal detection, we suppose the transmitted data is generated from 16-QAM modulation symbol.
IV-B Performance Evaluation for Channel Prediction
IV-B1 Baselines and Performance Metric
@@
  • Transformer: In [5], a transformer-based parallel channel predictor is proposed scheme to predict future channels in parallel, and thus avoid error propagation problems.
@@
  • LSTM: Long short-term memory (LSTM) [28] is designed with memory cells and multiplicative gates to deal with long-term dependency. An LSTM-based predictor with 4 LSTM layers is applied.
  • GRU: Gate recurrent unit (GRU) [29] is a variant of LSTM to address the vanishing gradient problem. Similarly, the number of GRU layers is set as 4 for channel prediction.
  • LLM4CP: [16] first attempts to propose a pre-trained LLM-empowered channel prediction method (LLM4CP) based on GPT-2 to achieve improved predictive capability and generalization ability.
  • Single-task LLM: We also train the proposed network on a single task to better analyze the performance.
For channel prediction task, NMSE is a widely used performance metric to directly characterize the accuracy of channel prediction. Therefore, NMSE is adopted as a major metric in our experiments.
IV-B2 Performance Analysis
@@
The testing dataset of channel prediction contains 10 velocities ranging from $10$ km/h to $100$ km/h, with $1000$ samples for each velocity. Fig. 4 reveals the NMSE performance of the proposed multi-task LLM with other baselines among different user velocities. It can be observed that as the user velocity increases, the NMSE of all baselines gradually increases. This is due to that as the user velocity increases, the CSI changes rapidly with a shorter channel coherence time, resulting in increased prediction difficulty. As illustrated in Fig. 4, attention-based methods, achieve relatively high-performance than traditional AI methods, validating the potential of attention-based method in channel prediction task. Equipped with excellent modeling capability of LLM, the proposed model finetuned on channel prediction task only, consistently outperforms other baselines among testing velocities. With elaborately designed multi-task instruction, the proposed LLM-enabled multi-task model obtains comparable channel prediction accuracy with single-task one, which verified the effectiveness of the proposed multi-task LLM.
IV-C Performance Evaluation for Multi-user Precoding
IV-C1 Baselines and Performance Metric
@@
  • ZF: A low complexity baseline is eigen-based zero-forcing (ZF) algorithm [30], which obtains the precoding matrix by the Moore-Penrose pseudo-inverse of the channel.
  • WMMSE: As mentioned above, the WMMSE algorithm [31] is one of the most popular iterative algorithms. The method can achieve near optimal sum rate performance while suffers from high computational complexity. The iterative number is set as $20$.
@@ IV-C2 Performance Analysis
The sum rate performance against different BS transmit power is plotted in Fig. 5. The noise power is set as $\sigma^{2}=-20$ dBW, the maximal transmit power of BS $P_{\rm max}$ increases from $-10$ to $20$ dBW. The ZF-based method and CNN-based method, though with low complexity, achieves unsatisfactory performance. The iterative algorithm WMMSE improves the sum rate, while it is still possible to fall in local optimal solutions. As depicted in Fig. 5, the proposed model, both trained on single task and trained on multiple tasks, achieves near-optimal performance for all the transmit powers, and outperforms other baselines.
IV-D Performance Evaluation for Signal Detection
IV-D1 Baselines and Performance Metric
@@
  • LMMSE: Linear minimum mean-squared error (LMMSE) detectors are a classical method for achieving signal detection with low complexity.
@@
  • OAMP-Net: A famous model-driven deep learning network proposed in [6], which incorporates deep learning into the orthogonal AMP (OAMP) algorithm.
@@
NMSE loss is the direct performance metric to present the accuracy of data recovery. It should be noted that the reason for choosing the data recovery accuracy rather than the bit error ratio (BER) or symbol error ratio (SER) to optimize and evaluate the model is described below. In the practical communication link, the objective of signal detection is to recover the data as accurately as possible, which can be utilized to compute the log-likelihood ratio (LLR) for demodulation and channel decoding to enable reliable communications. Thus, we utilized the NMSE to evaluate the performance of signal detection.
IV-D2 Performance Analysis
@@
Figure 6: The NMSE performance of proposed method and other baselines versus different SNRs.
@@
In Fig. 6, the NMSE performance of the proposed multi-task LLM on signal detection is presented. We also compare the performance with several baselines. The OAMPNet [6] achieves the best performance since it fully utilizes the statistical information of noise, which may not be obtained in practical system. Moreover, the proposed multi-task LLM can accurately recover the transmitted data, especially in low SNR regime. It can be observed that the performance of multi-task LLM is comparable to that of single-task LLM, suggesting the potential of multi-task LLM. Besides, the proposed method achieves high robustness in low SNR regime.
V Conclusions

### v2

@@ IV-C Multi-task Loss Function and Training Schedule
During the training stage, we randomly choose a task and sample data from the corresponding dataset for each iteration. After determining the involved task and data, we select the proper prompt and activate the required modules of encoder and decoder. It is worth noting that the backbone of the pre-trained LLM is frozen, while the other parameters of the network, together with the LoRA adapters are trainable. Then the proposed network update the parameters of encoders, decoders and LoRA adapters using the corresponding loss. The detailed loss functions for each task are illustrated below.
@@
For multi-user precoding, the training is divided into two stages, namely supervised learning and unsupervised learning, respectively. The WMMSE algorithm is one of the most popular iterative algorithms to find the locally optimal solutions. Therefore, in the supervised learning stage, we can first generate the power allocation vectors $\underline{\mathbf{p}}$ and $\underline{\mathbf{\lambda}}$ using the WMMSE algorithm as the label. Then, the supervised learning will employ the MSE loss function to make the power allocation vectors, $\hat{\mathbf{p}}$ and $\hat{\mathbf{\lambda}}$, generated by LLM as close to $\underline{\mathbf{p}}$ and $\underline{\bm{\lambda}}$ as possible, i.e.,
@@
$\text{Loss}_{Pre}=\frac{1}{2K}\left(\|\underline{\mathbf{p}}-\hat{\mathbf{p}}\|_{2}^{2}+\|\underline{\bm{\lambda}}-\hat{\bm{\lambda}}\|_{2}^{2}\right).$ (36)
Nevertheless, the WMMSE algorithm achieves only local optimality, making (36) insufficient for fully addressing the fundamental objective of problem P3. To enhance the overall rate performance, we implement additional network training using an unsupervised learning manner, where the loss function is directly derived from the original optimization objective:
@@
In the signal detection task, the original transmitted data is rearranged as $\mathbf{X}_{\rm Det}\in\mathbb{R}^{K\times 2}$, which serves as ground truth of the network output. Then we choose MSE loss for the training,
$\text{Loss}_{Det}=\frac{1}{2K}\|\mathbf{X}_{\rm Det}-\mathbf{X}_{\rm Det}^{O}\|_{F}^{2}.$ (38)
@@
V Simulation Results
In this section, extensive numerical simulations are presented to verify the effectiveness of the proposed method. Firstly, we elaborate on the simulation setup. Then the performance of three selected tasks, i.e. channel prediction, multi-user precoding, and signal detection, is evaluated respectively. Besides, the performance of LoRA fine-tuning-aware quantization is compared with the full-precision model. Finally, we analyze the impact of designed multi-task instruction and different LLM backbones, respectively.
V-A Simulation Setup and Training Details
For the experimental setup, we utilize the QuaDRiGa channel generator [28], implementing the 3GPP Urban Macro (UMa) propagation model [29] under non-line-of-sight (NLOS) conditions. The channel consists of 21 scattering clusters, each containing 20 propagation paths. We consider a multi-user MISO-OFDM system, where a BS simultaneously serves $K=4\sim 8$ moving users. BS employs a UPA comprising $N_{h}=16$ elements in horizontal and $N_{v}=8$ in vertical, while users are configured with single-antenna receivers. The antenna spacing is maintained at half of the wavelength at the center frequency. The users are uniformly distributed within angle range $[-\pi/2,\pi/2]$, and distance range $[\rho_{\rm min},\rho_{\rm max}]=[20~{}{\rm m},100~{}{\rm m}]$. We suppose a time-division duplex (TDD) system, where the center frequency of channel is set as $2.4$ GHz. The bandwidth of the channel is 8.64 MHz, comprising of $M=48$ subcarriers, i.e., the frequency interval of subcarriers is $180$ kHz. The dataset is partitioned into training and testing subsets, containing 50,000 and 10,000 samples per task respectively. The model undergoes training for 200 epochs over the dataset.
For channel prediction problem, We predict future CSI of $T_{2}=4$ time slots based on historical CSI of $T_{1}=16$ time slots. We suppose that the users are initialized with random positions and follow linear movement patterns. The velocity distribution for mobile users spans uniformly from $10$ km/h to $100$ km/h. To enhance the robustness against noise, the SNR is uniformly sampled between 5 dB and 20 dB during the training stage, to account for both the low-noise and high-noise scenarios. For uplink signal detection, we suppose the transmitted data is generated from 16-QAM modulation symbol. Similarly, the received signal is corrupted by noise with SNR uniformly distributed between 5 and 20 dB, during the fine-tuning.
V-B Performance Evaluation for Channel Prediction
V-B1 Baselines and Performance Metric
@@
  • Transformer: [5] introduces a parallelized channel prediction framework based on transformer to predict future CSI in parallel, and thus avoid error propagation problems.
@@
  • LSTM: The long short-term memory network (LSTM) [31] incorporates specialized memory units and gating mechanisms to effectively capture long-range temporal dependencies. Our implementation utilizes a four-layer LSTM structure for predictive modeling.
  • GRU: As an enhanced version of LSTM, the gated recurrent unit (GRU) [32] introduces simplified gating operations to mitigate gradient-related challenges during training. Similarly, the GRU-based model consists of 4 layers.
  • LLM4CP: LLM4CP  [13] represents a pioneering effort in leveraging pre-trained language model, i.e. GPT-2, for channel prediction task, demonstrating superior prediction accuracy and enhanced generalization.
  • Single-task LLM: We also train the proposed network on a single task to better compare the performance with the proposed multi-task LLM.
In channel prediction evaluation, the NMSE serves as a crucial indicator for assessing prediction precision, making it an essential measurement criterion in our experimental analysis.
V-B2 Performance Analysis
@@
The evaluation dataset for channel prediction comprises 10 distinct velocities spanning from $10$ km/h to $100$ km/h, with each velocity containing $1000$ data samples. As illustrated in Fig. 5, the NMSE performance of our proposed multi-task LLM framework is compared against various baseline methods across different user velocities. The historical CSI data is added by Gaussian whiten noise with ${\rm SNR}=20$ dB. Experimental results demonstrate a consistent degradation in NMSE performance across all methods as user mobility increases. This phenomenon can be attributed to the accelerated channel variation and reduced coherence time associated with higher velocities, which consequently amplifies the complexity of accurate channel estimation. Fig. 5 reveals that attention-based methods, achieve relatively high performance than traditional AI methods, validating the potential of attention-based methods in channel prediction task. Equipped with excellent modeling capability of LLM, the proposed model finetuned on channel prediction task only, consistently outperforms other baselines among testing velocities. With elaborately designed multi-task instruction, the proposed LLM-enabled multi-task model obtains comparable channel prediction accuracy with the single-task one, which verified the effectiveness of the proposed multi-task LLM.
Figure 6: The NMSE performance of proposed method and other baselines versus different SNR.
In Fig. 6, the robustness against noise of the proposed method is evaluated, where the SNR of noise in historical CSI is growing from 0 dB to 25 dB. The NMSE performance has been averaged over all test speeds. It can be observed that for all schemes, increased SNR conditions lead to improved NMSE performance in prediction accuracy. Thanks to the generalization ability of LLMs, the proposed method exhibits high robustness performance against CSI noise. It achieves the lowest NMSE performance in the entire SNR regime.
V-C Performance Evaluation for Multi-user Precoding
V-C1 Baselines and Performance Metric
@@
  • ZF: Eigen-based zero-forcing (ZF) algorithm [33] is a computationally efficient approach, which derives the precoding matrix through the Moore-Penrose pseudo-inverse operation applied to the channel matrix.
  • WMMSE: As mentioned above, the WMMSE algorithm [34] is one of the most popular iterative algorithms. The method can achieve satisfactory sum rate performance while suffering from high computational complexity. The iterative number is set as $20$.
@@ V-C2 Performance Analysis
The sum rate performance against different BS transmit power is plotted in Fig. 7. The noise power is set as $\sigma^{2}=-10$ dBW, the maximal transmit power of BS $P_{\rm max}$ increases from $-10$ to $10$ dBW. Besides, the user number is set as 4. As illustrated in Fig. 7, with the increase of the BS transmit power, the sum rate of all methods increase accordingly. As depicted in Fig. 7, the ZF-based method, though with low complexity, achieves unsatisfactory performance. The iterative algorithm WMMSE improves the sum rate, while it is still possible to fall in local optimal solutions, inducing an obvious gap from the capacity. With the powerful network and carefully designed training strategy, the deep-learning-based methods, including the CNN-based method and LLM-based method, are promising in conquering the problem and further improving the performance. Specifically, the proposed model, both trained on a single task and trained on multiple tasks, achieves near-optimal performance for all transmit power and outperforms other benchmark schemes for the entire transmit power range. Owing to the increasing size of the network, the LLM-based method exhibits superior optimization and generalization capabilities and outperforms the CNN-based method in terms of performance.
Figure 8: The sum rate performance of proposed method and other baselines versus different user numbers.
In Fig. 8, we illustrate the sum rate performance against the user number, which ranges from 4 to 8. The transmit power and noise power is set as 0 dBW and -10 dBW, respectively. As the user number increases, the spectrum efficiency increases with further exploitation of multiplexing gain. We observe from Fig. 8 that the proposed multi-task LLM-based method enjoys a higher sum rate performance, compared to existing methods for different number of users. This verifies the effectiveness and scalability of the proposed method. It is worth noting that LLMs inherently possess the ability to process variable-length sequences, thus our proposed scheme can be directly applied to different number of users without requiring any modifications. In contrast, for traditional AI methods, to accommodate varying numbers of users, input data under different user numbers must be zero-padded to match the shape of the maximum user number.
V-D Performance Evaluation for Signal Detection
V-D1 Baselines and Performance Metric
@@
  • LMMSE: Linear minimum mean-squared error (LMMSE) detector is a classical method for achieving signal detection with low complexity.
@@
  • OAMP-Net: A famous model-driven deep learning network proposed in [6], which incorporates deep learning into the orthogonal AMP (OAMP) algorithm for accurate signal detection.
@@
In this work, we utilize NMSE and symbol error ratio (SER) as metrics to evaluate the performance of signal detection. NMSE loss is the direct performance metric to present the accuracy of data recovery. It should be noted that the reason for employing the data recovery accuracy rather than only choosing the bit error ratio (BER) or SER to optimize and evaluate the model is described below. In the practical communication link, the objective of signal detection is to recover the data as accurately as possible, which can be utilized to compute the log-likelihood ratio (LLR) for demodulation and channel decoding to enable reliable communications. Besides, obtaining the recovered signal, we also demodulate the transmitted symbol by minimum euclidean distance decision, and then SER is employed to evaluate the performance in this case.
V-D2 Performance Analysis
@@
Figure 9: The NMSE performance of the proposed method and other baselines versus different SNRs.
@@
In Fig. 9, the NMSE performance of the proposed multi-task LLM on signal detection under different SNRs is presented. We also compare the performance with several baselines. As illustrated in Fig. 9, the DNN-based method presents poor performance since it directly inputs the data to the black-box-based network, while ignores the domain knowledge and structure of the problem and data. Besides, despite the low computational complexity, the LMMSE-based method also shows relatively high NMSE performance. The model-based methods, including DetNet [37] and OAMPNet [6], significantly improve the data recover accuracy; the OAMPNet [6] achieves higher performance than the LLM-based method in high SNR regime with better utilization of the statistical information of noise. Moreover, the proposed multi-task LLM can accurately recover the transmitted data, especially in the low SNR regime. It can be observed that the performance of multi-task LLM is comparable to that of single-task LLM, suggesting the potential of multi-task LLM networks in wireless communications.
The SER performance against SNR is depicted in Fig. 10. When the SNR is lower than 8 dB, the proposed multi-task LLM network outperforms other baselines, which indicates that the generalization capability of LLMs endows the proposed method with high robustness against noise. As the SNR increases, the OAMPNet [6] achieves the best performance, since it fully utilizes the statistical information of noise. Based on this observation, we provide two comments as follows. First, for the scenarios where the power of noise $\sigma^{2}$ can be obtained, future works can consider effectively incorporating statistical information of channel and noise as prompt to further improve the performance of LLM-based method, especially in high SNR regimes. Secondly, there are still many cases where the statistical information of noise may not be obtained in practical system. In these cases, OAMPNet may fail to achieve satisfactory performances.
Figure 10: The SER performance of the proposed method and other baselines versus different SNRs.
V-E Performance evaluation for LoRA-fine-tuning-aware quantization
For fair and convenient comparison, in the above subsections, we mainly utilize full-precision model for performance evaluation. Then in this subsection, we focus on the impact of LoRA-fine-tuning-aware quantization.
TABLE I: Comparison of the model with LoRA-fine-tuning-aware quantization with other baselines.
Quantization method NMSE for CP NMSE for DET Sum rate for PRE
Full-precision -16.37 dB 0.0322 29.3658 bit/s/Hz
LoRA-fine-tuning-aware quantization -16.25 dB 0.0363 29.3632 bit/s/Hz
Traditional quantization -15.68 dB 0.0528 28.6032 bit/s/Hz
$\bullet$ “CP” denotes channel prediction, “DET” denotes signal detection, and “PRE” denotes multi-user precoding.
TABLE II: Comparison of the model with different LLM backbone.
LLM backbone NMSE for CP NMSE for DET Sum rate for PRE Network parameters
LLAMA2-7B -16.37 dB 0.0322 29.3658 bit/s/Hz 23.9713 M
GPT2 -15.39 dB 0.0573 29.0365 bit/s/Hz 17.3338 M
- -13.49 dB 0.1908 27.2981 bit/s/Hz 19.777 M
$\bullet$ “CP” denotes channel prediction, “DET” denotes signal detection, and “PRE” denotes multi-user precoding.
The performance comparison of the proposed full-precision model, the model with LoRA-fine-tuning-aware quantization, and the model with traditional quantization are presented in Table I. For the full-precision model, the model parameters are stored and computed in a 16-bit floating-point format. On the contrary, the model with LoRA-fine-tuning-aware quantization transforms the LLM backbone into a 4-bit integer format, which indicates a 75% storage reduction during the fine-tuning. Besides, the iterative number for initialization is set as $Iter=5$. It is noted here that during LoRA fine-tuning, the quantized weight is temporarily dequantized to the simulated high-precision weight to facilitate accurate computation. Moreover, the model with traditional quantization serves as a baseline, which quantifies the LLM backbone weight matrices, initials the adapters as (29), and conducts LoRA fine-tuning afterward.
As shown in Table I, the model with LoRA-fine-tuning-aware quantization achieves comparable performance with the proposed full-precision model. For the channel prediction task, the NMSE is averaged for different speeds and SNRs; for the signal detection task, the performance is averaged for different SNRs; for multi-user precoding task, we set user number as 4, the power of noise and transmit power is set as -10 dBW and 0 dBW, respectively. Besides, it significantly outperforms the model with traditional quantization for all the tasks selected. We can conclude from the results that the introduced LoRA-fine-tuning-aware quantization successfully approximates high-precision weight by the quantized weight and low-rank adapters, and thus the performance degradation resulting from quantization is negligible.
V-F Impact of multi-task instruction
In this subsection, we analyze the influence of the proposed multi-task instruction module in Section III-A on the performance and convergence rate of neural network training. It is worth noting that, in order to distinguish different tasks, utilizing multi-task instruction as prompts is indispensable for the proposed multi-task PHY network to understand different task requirements. For single-task LLM, although it is possible to neglect the prompt, introducing the designed instruction with wireless data can involve domain knowledge to facilitate task-specific adaptation of LLMs.
Figure 11: Training losses of different methods against epoch.
In Fig. 11, we take the signal detection task for instance to illustrate the training loss in MSE against the training epoch. In this figure, we depict the training loss of the proposed model trained for multiple tasks with prompts, the proposed model trained for a single task with prompts, and the proposed model trained for a single task without prompts. We can observe that the overall trends of all losses are declining with the increase of training epoch. The training loss of single-task LLM with prompts converges the fastest, and 200 epochs of training are enough to achieve satisfactory performance. When the training epoch reaches about 200, the loss function of multi-task LLM also converges. Nevertheless, the proposed model trained for a single task without prompts requires 300 epochs of training to obtain comparable performance with the model with prompts. Therefore, it is indicated that the incorporation of designed multi-task instruction as prompts significantly accelerates the network fine-tuning and improves the LLM’s adaptability to downstream tasks.
V-G Impact of LLM backbone
Finally, in this subsection, the influence of LLM backbone on the performance is evaluated. In this work, we employ LLAMA2-7B as the backbone. To analyze the impact, we compare the selected backbone with the following benchmarks. Firstly, we replace the LLAMA2-7B with a smaller pre-trained model GPT2 [8] while other modules remain the same. Secondly, the LLM backbone is directly removed, which means that the output of the encoders is straightly fed into the designed decoders.
We summarize the simulation results in Table II. Similarly, as stated in Section V-E, the NMSE is averaged within different SNRs and speeds for channel prediction and is averaged within different SNRs for signal detection. Besides, user number is set as 4, the power of noise and transmit power is set as -10 dBW and 0 dBW for multi-user precoding. It can be observed from Table II that the performance achieved in different tasks improves with the introduction of the LLM backbone. For instance, the NMSE performance for channel prediction of the proposed method with LLAMA2 backbone is -16.37 dB, while the performance drops 3 dB when removing the backbone. And the performance drops sharply especially for signal detection task. Furthermore, due to the increasing size of the LLM backbone, the model with LLAMA2-7B backbone outperforms the model with GPT2 backbone, although GPT2 can still achieve the multi-task PHY network. Therefore, we can adopt a proper LLM backbone to balance the computational costs and performance.
VI Conclusions